BG February Com Night 6, UT February 6 2009

Observers: J.Hill, A.Rakich, N.Ageorges, V.Kneirim (in Tucson) plus J.Storm (in Potsdam)
Telescope Operator: D. Gonzalez Huerta
SW Support: M.Delapena (on-call)
Telescope Support: J. Rousseau

Summary:

Closed for high winds most of the night. Only got a few minor things tested.

Preparations:

Details (Times in MST):

Open

18:05 Open Dome, T=+3.5 degC. D=-11 degC, Wind (LBTO) high.

Having a problem that the LUCIFER V880 Sun machine is not talking on the network.

18:25 pointto M5_1133 gs=0 ACQUIRE PAM- Adjust IE=-18 CA=8

18:35 pointto BS9122 gs=0 ACTIVE - Initial seeing is 1.2 arcsec at EL=47 deg

Close

18:42 Closed for high wind.

19:04 Editted lbt.conf to set default GCS.guideRate to 0.5, and default GCS.cyclesUntilStable to 1.

19:21 Closed dome preset sent by LUCIFER. TRACK

M3 tests for pupil alignment

19:30 Park telescope at zenith for M3 tests.

19:42 M3 tip from -615" to 0". Sending the update from PSF does nothing. See Issue 1997.

19:44 Adjusted M3 tip from OSS directly. -615, 0 615, -615, -915, -1215. This confirms that -615 tilt is clearly better alignment than +615, so that bit of historical mystery is resolved.

20:03 Adjusted M3 tilt from OSS directly -20, -320,

20:20 Measuring pupil at different elevation angles. 90, 60, 30, 0.

Polycom communication to the to-station is not working. Both units can dial other places, but they can't dial each other. ??? Power cycling both of them doesn't help.

20:46 focus 0 to test Issue #1851 a bit.

Open

21:39 Reopen as the wind is decreasing and it seems we have a couple hours of clear sky.

21:45 pointto BS9122 gs=0 ACTIVE - Why was M2 offload forcing M1 into the -X and -Z limits? Was it the stepFocus commands I issued while the telescope was parked?

21:47 pointto BS9122 gs=0 ACTIVE - succeeded after we cleared the large M2 offload.

21:51 stray light in these WFS images. Probably much of it is coming from moon-illuminated clouds, but we don't understand the sharp top-bottom pattern.

21:58 offset 10 0 - took 42 sec.

Close

22:12 Closing again

22:30 Jesper joins us from Potsdam, and we get Polycom communication by all dialing into the unit in Room N505.

Open

01:45 Reopen the dome

01:50 pointto BS9122 gs=0 ACTIVE - Why was M2 offload forcing M1 into the -X and -Z limits? Was it the stepFocus commands I issued while the telescope was parked?

01:54 pointto BS9122 gs=0 ACTIVE - still failed because star is far from focus

01:56 pointto BS9122 gs=0 ACQUIRE PAM-

01:58 pointto BS9122 gs=0 ACTIVE - succeeded - Seeing is 1.6 arcsec at EL=45

02:05 pointto BS9122 gs=1 ACTIVE - couldn't move probe

02:06 pointto BS9122 gs=1 PA=0 ACTIVE - couldn't move probe

02:07 pointto BS9122 gs=25 PA=0 ACTIVE - problems with M1 collimation

02:08 pointto BS9122 gs=25 PA=0 ACTIVE

02:13 pointto BS9122 gs=31 PA=0 ACTIVE - no star found (too faint)

02:16 pointto BS9122 gs=21 PA=0 ACTIVE - no star found (too faint)

02:17 pointto BS9122 gs=17 PA=0 ACTIVE - no star found (too faint)

02:18 pointto BS9122 gs=6 PA=0 ACTIVE -- guiding was not stable within 6 cycles ??

02:20 pointto BS9122 gs=6 PA=0 ACTIVE -- guiding was not stable within 6 cycles ??

02:22 pointto BS9122 gs=0 PA=0 ACTIVE -- succeeds on-axis

02:24 pointto BS9122 gs=6 PA=0 ACTIVE -- guiding was not stable within 6 cycles ??

02:26 pointto BS9122 gs=44 PA=0 ACTIVE -- now I have seen one of the spiraling in cases, and captured the time! AZ=-51:40 EL=40:49 See Issue #2000.

02:30 pointto BS9122 gs=36 PA=0 ACTIVE - no star found (too faint)

countThreshold=50, threshMultiplier=3.0 (I adjusted both quickly because I'm not certain which mode we are in.)

02:35 pointto BS9122 gs=36 PA=0 ACTIVE - succeeded but hampered by clouds.

02:36 Close for wind

Close

HH:MM Close dome T=?degC, D=?degC, wind (LBTO) ? m/sec.

Planned Activities

  • LUCIFER tests with Build 20 (done in closed dome)

  • Pupil measurements versus M3 tip-tilt (done in closed dome)

  • Confirm start of guiding with waiting only 1 cycle (waiting for sky, we saw it work 1 time)

  • Measure more coordinates for Jesper's AGW transformations (highest priority).

  • Measure offset timings

-- JoarBrynnel - 15 Jan 2009
Topic revision: r5 - 06 Feb 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback